OpenStudy (anonymous):

find the solutions of the system y=x^2+3x-4 y=2x+2 I would prefer if you could walk me through the question rather than just give me the answer, but whatever you can do for me is appreciated

OpenStudy (anonymous):

well do you know where to begin or have no idea

OpenStudy (anonymous):

I was absent a couple of days this is a review I honestly have no clue what to do, I hate getting sick

OpenStudy (anonymous):

lol no problem. well with any system the main objective is to find an x,y, etc. value that satisfies both equations

OpenStudy (anonymous):

It's multiple choice if that helps at all, it seems I need two pairs of (x,y)

OpenStudy (anonymous):

you can probably start by letting 2x+2=x^2+3x-4

OpenStudy (anonymous):

So, I combine the two problems?

OpenStudy (anonymous):

Then I need to find an (x,y) that will work for both?

OpenStudy (anonymous):

Well, by setting the two functions equal to each other you can solve for the two values of x and plug each back into the equation to find the two corresponding y values

OpenStudy (anonymous):

So I need to start by finding the two x's? alright that should be pretty easy

OpenStudy (anonymous):

yep, just factor and find the x values, then plug back in to find the y values.

OpenStudy (anonymous):

I feel like there's no solution, but I don't want to say that, because the only reason I came to that conclusion is because I looked at the possible answers and none of them fit

OpenStudy (anonymous):

what are the possible answers

OpenStudy (anonymous):

A. (-3,6) and (2,-4) B.(-3,-4) and (2,6) C.(-3,-4) and (-2,-2) D. No solution

OpenStudy (anonymous):

because, either way -3 is in there, but neither -2 or 2 make them equal

OpenStudy (anonymous):

the answer is b

OpenStudy (anonymous):

But, the x's don't equal out..

OpenStudy (anonymous):

oh, they don't have to equal. the sole purpose of setting the two equal to each other was to solve for the x values. once you have them, you plug back into either equation to solve for the y values

OpenStudy (anonymous):

wait, so how would I know which x values go in there? Really sorry about this but I would greatly appreciate it, if you could walk me through all the steps, I'm definetly going to give you a medal

OpenStudy (anonymous):

the x values you obtained satisfy both equations, so you can plug them into either one of the original functions to obtain the corresponding y values

OpenStudy (anonymous):

Wait, how does it satisfy the equation, I'm so confused right now ;_;

OpenStudy (anonymous):

So what I mean by satisfy the equation is this: a value for x that will make the equation true. so for example: after setting the two functions equal to each other and some algebra we get this: x^+x-6=0 so looking at this equation we see that it is equal to 0, so the question is: what values for x will make this equation true or what values will make the expression sum to 0

OpenStudy (anonymous):

okay, if we have x^2+x-6=0 we need to make 0 with the 2 x factors? how does B make that 0 it would be -3^ = -9+2-6 that doesn't equal 0 and 2^2 = 4-3-6 also doesn't equal 0

OpenStudy (anonymous):

unless we sub in y for the 0

OpenStudy (anonymous):

so we factor and get x=2,-3 next we want to find our y values. well plugging these x values into to either of the original function will give us the two y values. y=2x+2 ==> plug in x=2 y=2(2)+2=6 set 1= (2,6) y=2x+2 ==> plug in x=-3 y=2(-3)+2=-4 set 2= (-3,-4)

OpenStudy (anonymous):

The x values you found makes x^2+x-6=0 B is the set of x and y values that satisfy your original system of equation. Here's how you check: lets take set 1=(2,6) we can plug it in to either equation. Ill choose the easier one. y=2x+2 ===> plug in set 1 6=2(2)+2 this set of numbers satisfies the equation does that make sense?

OpenStudy (anonymous):

I understand that it works in the main problem, but what I don't understand is how you said that the x values work for the other problem

OpenStudy (anonymous):

what other problem? you mean x^2+x-6=0 ?

OpenStudy (anonymous):

So what I mean by satisfy the equation is this: a value for x that will make the equation true. so for example: after setting the two functions equal to each other and some algebra we get this: x^+x-6=0 so looking at this equation we see that it is equal to 0, so the question is: what values for x will make this equation true or what values will make the expression sum to 0

OpenStudy (anonymous):

You said that earlier how do the values for x make it = 0? unless you're not supposed to sub them in in that problem

OpenStudy (anonymous):

well, x^2-3x-1=8x-1

OpenStudy (anonymous):

that turns into x^2-11x=0

OpenStudy (anonymous):

Now that, that equals 0, we need to factor and we get (0,11). Right?

OpenStudy (anonymous):

So, now we put that into the main problems to find the y y=0^2-0-1 y=-1 y=8(11)-1 88-1 y=87, so it's be answer b again (0,-1) and (11,87) right?

OpenStudy (anonymous):

If so, then I understand it, and thank you

OpenStudy (anonymous):

why do you think we set the two equations equal to each other?

OpenStudy (anonymous):

To solve for the x values?

OpenStudy (anonymous):

Well, either way number 11 was in the back of the book and it was answer b, so hank you I got it now I gave you a metal

OpenStudy (anonymous):

yes, to solve for the x values that make x^2+x-6=0 now, x^2+x-6=(x+3)(x-2)=0 from here we can see that x=-3,2 satisfies the equation but it is not so obvious for X^2+x-6=0 which is why you always have to factor polynomial equations

OpenStudy (anonymous):

Omg, open study updates so slowly, I'm sorry but I closed the question

OpenStudy (anonymous):

okay well goodluck

OpenStudy (anonymous):

thank you, for the luck and the help.